# Off-policy Evaluation with Deeply-abstracted States

### v2

@@ D.4 Proof of Theorem 2
Additionally, at the first iteration, these Fisher consistencies directly follows from Lemma 1. Consequently, it suffices to prove the Fisher consistencies at later iterations. Below, we first prove the Fisher consistencies of SIS, MIS and Q-function-based estimator at the second iteration. Next, we prove the resulting abstraction is a Markov state abstraction (Allen et al.,, 2021) in that the data generating process when confined to the abstract state space remains an MDP. This together with Lemma 1 proves the Fisher consistencies at the third iteration. Using similar arguments, we can establish the Fisher consistencies at any $K>3$ iterations. The proof can thus be completed.
D.4.1 Fisher consistencies at the 2nd iteration
It is worthwhile mentioning that at the second iteration, the refined backward-model-irrelevance condition is defined with respect to the abstract state space induced by the forward abstraction $\phi_{1}$ at the first iteration instead of the ground state space. In particular, we require
$b_{\phi_{1},t}(a_{t}|x_{t}^{(1)},a_{t-1},x_{t-1}^{(1)},\cdots,x_{1}^{(1)})=b_{\phi_{1},t}(a_{t}|x_{t}^{(2)},a_{t-1},x_{t-1}^{(2)},\cdots,x_{1}^{(2)}),$ (D.11)
for any $t$ and $\{a_{t}\}_{t}$, whenever $\phi_{2}(x_{t}^{(1)})=\phi_{2}(x_{t}^{(1)})$ for any $\{x_{t}^{(1)}\}_{t}$ and $\{x_{t}^{(2)}\}_{t}$, where $b_{\phi_{1},t}$ denote the history-dependent behavior policy (see also the denominator of Equation D.6), and
$\begin{split}&\sum_{x\in\phi_{2}^{-1}(x_{2})}\mathbb{P}(A_{t}=a,\phi_{1}(S_{t})=x|\phi_{1}(S_{t+1})=x^{(1)})\\ =&\sum_{x\in\phi_{2}^{-1}(x_{2})}\mathbb{P}(A_{t}=a,\phi_{1}(S_{t})=x|\phi_{1}(S_{t+1})=x^{(2)}),\end{split}$ (D.12)
whenever $\phi_{2}(x^{(1)})=\phi_{2}(x^{(2)})$.
  • Fisher consistency of SIS. When restricting to the abstract state space induced by $\phi_{1}$, the resulting behavior policy is not guaranteed to be Markovian. To address this challenge, SIS employs the history-dependent IS ratio defined in Equation D.6 to maintain consistency. Let $\rho_{1:t,\phi_{1}}^{\pi}$ and $\rho_{1:t,\phi_{2}\circ\phi_{1}}^{\pi}$ denote the history-dependent SIS ratios at the first and second iterations, respectively. Under $\pi$-irrelevance and the refined history-dependent-behavior-policy-irrelevance (see (D.11)), it is immediate to see that $\rho_{1:t,\phi_{1}}^{\pi}=\rho_{1:t,\phi_{2}\circ\phi_{1}}^{\pi}$ so that $\rho^{\pi}$-irrelevance is achieved at the second iteration. This in turn validates the unbiasedness of the SIS estimator based on $\{\rho_{1:t,\phi_{2}\circ\phi_{1}}^{\pi}\}_{t}$.
  • Fisher consistency of MIS. We first show that the abstraction produced by DSA at the second iteration achieves $w^{\pi}$-irrelevance, i.e., $w^{\pi}_{\phi_{1}}=w^{\pi}_{\phi_{2}\circ\phi_{1}}$. We next establish the identifiability of the MIS ratio $w^{\pi}_{\phi_{2}\circ\phi_{1}}$.
    The proof is very similar to that of Theorem 1. Specifically, define
    $w_{t,\phi_{1}}^{\pi}(a,x)=\frac{\mathbb{P}^{\pi}(A_{t}=a,\phi(S_{t})=x)}{\mathbb{P}(A_{t}=a,\phi(S_{t})=x)},$
    When $t=1$, $w_{1,\phi_{1}}^{\pi}$ is reduced to the IS ratio $\pi_{\phi_{1}}(a|x)/b_{1,\phi_{1}}(a,x)$. Under $\pi$-irrelevance and behavior-policy-irrelevance (by setting $j$ in Equation D.11 to $1$), the numerator $\pi_{\phi_{1}}$ and denominator $b_{1,\phi_{1}}$ equal $\pi_{\phi_{2}\circ\phi_{1}}$ and $b_{1,\phi_{2}\circ\phi_{1}}$ (the behavior policy when restricting to the abstract state space produced by DSA at the 2nd iteration), respectively. This establishes the irrelevance in $w_{1,\phi_{1}}^{\pi}$.
    $\begin{split}&\mathbb{P}(A_{t}=a_{2}|\phi_{1}(S_{t})=x_{2}^{(1)},A_{t-1}=a_{1},\phi_{1}(S_{t-1})=x_{1}^{(1)})\\ =&\mathbb{P}(A_{t}=a_{2}|\phi_{1}(S_{t})=x_{2}^{(2)},A_{t-1}=a_{1},\phi_{1}(S_{t-1})=x_{1}^{(2)}),\end{split}$ (D.13)
    Let $X_{t}$ denote $\phi_{1}(S_{t})$ for any $t$. We next claim that
    $w_{t+1,\phi_{1}}(a,x)=\mathbb{E}\Big{[}w_{t,\phi_{1}}(A_{t},X_{t})\frac{\pi_{\phi_{1}}(A_{t}|X_{t})}{b_{2,\phi_{1}}(A_{t+1}|X_{t+1},A_{t},X_{t})}|A_{t+1}=a,X_{t+1}=x\Big{]}.$ (D.14)
    Notice that this formula is very similar to (D.9). The only difference lies in that the denominator of the IS ratio on the RHS is no longer Markovian. Rather, it depends on the current state as well as the previous state-action pair. Meanwhile, (D.14) can be proven using similar arguments to (D.9).
    Based on (D.14), we are ready to establish the irrelevance in $w_{t+1,\phi}^{\pi}$. In particular, looking at the RHS of (D.14), both $w_{t,\phi_{1}}$ and the IS ratio $\pi_{\phi_{1}}/b_{2,\phi_{1}}$ depend on $X_{t}$ and $X_{t+1}$ only through their abstractions $\phi_{2}(X_{t})$ and $\phi_{2}(X_{t+1})$. Meanwhile, the conditional distribution of $A_{t},X_{t}$ given $A_{t+1},X_{t+1}$ depends on $X_{t}$ and $X_{t+1}$ through their abstractions, as well, given (D.13) and (D.12). This establishes the irrelevance in $w_{t+1,\phi_{1}}$. By induction, we have proven the irrelevance in $w_{t,\phi_{1}}$ for any $t$. Under the coverage assumption in Assumption 2, these ratios are uniformly bounded. It follows that the limit $\lim_{T}\sum_{t=1}^{T}\gamma^{t-1}w_{t,\phi_{1}}^{\pi}$ is well-defined. By setting $T\to\infty$, we obtain the irrelevance in $w_{\phi_{1}}^{\pi}$.
    So far, we have established the $w^{\pi}$-irrelevance. This in turn yields $\mathbb{E}[f_{3}(w_{\phi_{1}}^{\pi})]=\mathbb{E}[f_{3}(w_{\phi_{2}\circ\phi_{1}}^{\pi})]$, according to Lemma D.1. It remains to prove the identifiability of $w_{\phi_{2}\circ\phi_{1}}^{\pi}$. However, this can be proven using similar arguments to the proof of Theorem 1. Specifically, we first observe that $w_{\phi_{2}\circ\phi_{1}}^{\pi}=\lim_{T}\sum_{t=1}^{T}\gamma^{t-1}w_{t,\phi_{2}\circ\phi_{1}}^{\pi}$. Next, when setting $t=1$, the identifiability of $w_{1,\phi_{2}\circ\phi_{1}}^{\pi}$ is readily available, given that of $\rho_{\phi_{2}\circ\phi_{1}}^{\pi}$. Finally, since $w_{t+1,\phi_{2}\circ\phi_{1}}(a,x_{2})$ equals
    we can employ similar arguments to the proof of Theorem 1 to prove the identifiability of the above expression, assuming $w_{t,\phi_{2}\circ\phi_{1}}$ is identifiable. By induction, this establishes the identifiability of $w_{\phi_{2}\circ\phi_{1}}$.
  • Fisher consistency of Q-function-based method. The Fisher consistency of Q-function-based method can be established in a similar manner to that in Theorem 1. Specifically, under $\pi$-irrelevance, it is trivial to show $\mathbb{E}[f_{1}(Q^{\pi}_{\phi_{1}})]=J(\pi)=\mathbb{E}[f_{1}(Q^{\pi}_{\phi_{2}\circ\phi_{1}})]$. Meanwhile, its identifiability is readily obtained based on the results in the following section, which proves that the process $(\phi_{2}(\phi_{1}(S_{t})),A_{t},R_{t})_{t\geq 1}$ remains an MDP.
D.4.2 Backward abstraction is a Markov state abstraction
@@
It is equivalent to prove that, when the backward abstraction $\phi$ is obtained by applying the refined backward-model-irrelevance condition to the original MDP $(S_{t},A_{t},R_{t})_{t\geq 1}$ with a history-dependent behavior policy, the reduced process $(\phi(S_{t}),A_{t},R_{t})_{t\geq 1}$ remains an MDP.
@@
Proof of Lemma D.2. Using similar arguments to (D.2), we have
@@
where the third equation is due to the backward-transition-irrelevance condition, under which $\mathbb{P}(A_{t}=a,\phi(S_{t})=x|S_{t+1}=s_{t+1})$ equals $\mathbb{P}(A_{t}=a,\phi(S_{t})=x|S_{t+1}=s_{t+1}^{\prime})$. This proves (D.15).
@@
Using similar arguments to (D.2) and (D.17), the LHS can be represented by
@@
Proof of the Markov property. We next prove that the refined backward abstraction is indeed an MSA, despite that the behavior policy is no longer Markovian. Toward that end, we first show that the evolution of $\phi(S_{t})$ remains Markovian. Specifically, we aim to show
@@
where the last equation again, follows from the backward-transition-irrelevance.
@@
Now, using the same arguments to (D.4.2) and (D.26), we obtain
@@
D.5 Lemma D.3 and its proof
We first state Lemma D.3.
Lemma D.3
@@
  • The forward abstraction selects the first two groups $S_{t}^{(1)}$ and $S_{t}^{(2)}$;
  • The proposed backward abstraction selects the last two groups $S_{t}^{(2)}$ and $S_{t}^{(3)}$;
@@
We next prove this lemma. Notice that reward-irrelevance requires the reward function (i.e., the conditional mean of the immediate reward given the state-action pair) to depend on the state only through its abstraction. Under the deterministic reward assumption in Lemma D.3, such a conditional mean independence is equivalent to conditional independence. In other words, reward-irrelevance is achieved if the reward is conditionally independent of the state given the action and the abstract state.
D.5.1 Proof for the bandit example

### v3

@@ D.4 Proof of Theorem 2
In this subsection, we will prove each part of Theorem 2 individually. Recall that $\phi$ denote an MSA from $\mathcal{S}$ to $\mathcal{X}$ and $\phi^{*}$ denote another MSA from $\mathcal{X}$ to certain $\mathcal{X}^{*}=\{\phi^{*}(x):x\in\mathcal{X}\}$. Then we have:
  • When lookup tables are employed to parameterize the Q-function and MIS ratio, under certain matrix invertibility condition detailed in Appendix D, we have
    $\kappa_{q}(\phi)=O\Big{(}\frac{{|\mathcal{X}||\mathcal{A}|}}{\sqrt{NT}}\Big{)}\,\,\hbox{and}\,\,\kappa_{w}(\phi)=O\Big{(}\frac{{|\mathcal{X}||\mathcal{A}|}}{\sqrt{N}}\Big{)}.$ (D.34)
    This result directly follows from Lemma D.2.
  • In this section, we first prove that for any MSA $\phi$, we have $\textrm{Var}^{(1)}\geq\textrm{Var}^{(1)}(\phi)$ with infinite horizon $T$, where $\textrm{Var}^{(1)}$ denotes the MSE of DRL applied to the original ground state space. It follows from (3) that, for any MSA $\phi$, the MSE of DRL is
    $\textrm{Var}^{(1)}(\phi)=$ $\frac{1}{N}\mathbb{V}\left[V_{\phi}^{\pi}\left(\phi(S)\right)\right]+\frac{1}{NT(1-\gamma)^{2}}\mathbb{V}\left[w_{\phi}^{\pi}(A,\phi(S))\left(R+\gamma V_{\phi}^{\pi}\left(\phi(S^{\prime})\right)-Q_{\phi}^{\pi}(A,\phi(S))\right)\right]$
    $+O\Big{(}\frac{R_{\max}^{2}\max(\kappa_{q}(\phi),\kappa_{w}(\phi),N^{-1/2})\log(N)}{(1-\gamma)^{2}N}+\frac{R_{\max}^{2}\max(\kappa_{q}(\phi)\kappa_{w}(\phi),\sqrt{N}\kappa_{q}^{2}(\phi)\kappa_{w}^{2}(\phi))}{\sqrt{N}(1-\gamma)^{2}}\Big{)}$
    Under an infinite horizon, we have
    $\textrm{Var}^{(1)}(\phi)=$ $\frac{1}{N}\mathbb{V}\left[V_{\phi}^{\pi}\left(\phi(S)\right)\right]$
    $+O\Big{(}\frac{R_{\max}^{2}\max(\kappa_{q}(\phi),\kappa_{w}(\phi),N^{-1/2})\log(N)}{(1-\gamma)^{2}N}+\frac{R_{\max}^{2}\max(\kappa_{q}(\phi)\kappa_{w}(\phi),\sqrt{N}\kappa_{q}^{2}(\phi)\kappa_{w}^{2}(\phi))}{\sqrt{N}(1-\gamma)^{2}}\Big{)}.$
[…]
    $\mathbb{V}\left[\mathbb{E}\left\{V^{\pi}\left(S\right)|\phi(S)\right\}\right]+\mathbb{E}\left[\mathbb{V}\left\{V^{\pi}(S)|\Phi(S)\right\}\right]=\mathbb{V}\left[V^{\pi}\left(S\right)\right],$
    we have
    $\mathbb{V}\left[V_{\phi}^{\pi}\left(\phi(S)\right)\right]+\mathbb{E}\left[\mathbb{V}\left\{V^{\pi}(S)|\Phi(S)\right\}\right]=\mathbb{V}\left[V^{\pi}\left(S\right)\right].$
    This means that after one Markov state abstraction, the MSE of DRL will decrease, that is, $\lim_{T\to\infty}\textrm{Var}^{(1)}\geq\lim_{T\to\infty}\textrm{Var}^{(1)}(\phi)$.
    Since both $\phi$ and $\phi^{*}$ are MSA, then $\phi^{*}\circ\phi_{1}$ is also an MSA. So we have $\lim_{T\to\infty}\textrm{Var}^{(1)}(\phi)\geq\lim_{T\to\infty}\textrm{Var}^{(1)}(\phi^{*}\circ\phi)$.
    Additionally, from Lemma D.2 and equation (D.34), we can conclude that $\kappa_{q}$ and $\kappa_{w}$ are proportional to $|\mathcal{S}|$, while $\kappa_{q}(\phi)$ and $\kappa_{w}(\phi)$ depend on $|\mathcal{\phi(S)}|$, meaning the bias term will also decrease.
  • We know that the conditional expectation $\mathbb{E}[\rho_{1:t}^{\pi}(A_{i,j},S_{i,j})|A_{i,j},\phi(S_{i,j})]=\rho_{\phi,1:t}^{\pi}(A_{i,j},\phi(S_{i,j}))$ holds.
    By applying Jensen’s inequality, we can derive the following result:
    $\mathbb{E}[(\rho_{1:j}^{\pi})^{2}(A_{i,j},S_{i,j})]\geq\mathbb{E}[(\rho_{\phi,1:j}^{\pi})^{2}(A_{i,j},\phi(S_{i,j}))]$
    This inequality implies that after performing one Markov state abstraction operation, the upper bound of the MSE for the SIS method will decrease. Hence, by deduction, we have $\textrm{Var}^{(2)}(\phi)\geq\textrm{Var}^{(2)}(\phi^{*}\circ\phi)$.
  • The bias term in (4) directly follows Assumption 1 and 2, which is independent of $\phi$.
D.5 Proof of Lemma 2
@@
We provide the more complex version, where the type-II MSA $\phi$ is obtained by applying the refined backward-model-irrelevance condition to the original MDP $(S_{t},A_{t},R_{t},S_{t+1})_{t\geq 1}$ with a history-dependent behavior policy. In this case, the reduced process $(\phi(S_{t}),A_{t},R_{t},\phi(S_{t+1}))_{t\geq 1}$ remains an MDP. The type-II MSA $\phi$, obtained by applying the original condition, directly follows from this refinement.
@@
Proof of Lemma D.3. Using similar arguments to (F.2), we have
@@
where the third equation is due to the backward-transition-irrelevance (7) condition, under which $\mathbb{P}(A_{t}=a,\phi(S_{t})=x|S_{t+1}=s_{t+1})$ equals $\mathbb{P}(A_{t}=a,\phi(S_{t})=x|S_{t+1}=s_{t+1}^{\prime})$. This proves (D.35).
@@
Using similar arguments to (F.2) and (D.37), the LHS can be represented by
@@
Proof of the Markov property. We next prove that the refined type-II abstraction is indeed an MSA, despite that the behavior policy is no longer Markovian. Toward that end, we first show that the evolution of $\phi(S_{t})$ remains Markovian. Specifically, we aim to show
@@
where the last equation again, follows from the backward-transition-irrelevance (7).
@@
Now, using the same arguments to (D.5) and (D.46), we obtain
@@
Appendix E Toy examples
To elaborate the usefulness of DSA in reducing state cardinality, we analyze two examples: a bandit example and an MDP example. In both examples, we focus on a specific type of state abstraction known as variable selection, which selects a sub-vector from the original state. Additionally, we focus on the class of state-agnostic target policies where $\pi$ is independent of the states. This type of policy is prevalent in causal inference and A/B testing, where the objective is to learn the global treatment effect of applying either a new or old policy consistently over time [see e.g., Hu and Wager, 2022, Leung, 2022, Bojinov et al., 2023, Shi et al., 2023, Xiong et al., 2024].
Figure E.1: Two illustrative examples.
Example I (A bandit example): To illustrate the main idea, we start by considering the contextual bandit setting [CB, Dudík et al., 2014], which can be regarded as a special MDP with independent state transitions. Under this setting, the states are i.i.d. generated, and model-irrelevance is reduced to reward-irrelevance whereas the proposed backward-model-irrelevance simplifies to behavior-policy-irrelevance. When specialized to variable selection in CB, our proposal is reduced to the iterative confounder selection algorithm in causal inference [see e.g., Guo et al., 2022]; see also the review of confounder selection in Appendix A. We assume the states can be divided into three independent groups, denoted by $S_{t}^{(1)}$, $S_{t}^{(2)}$ and $S_{t}^{(3)}$, respectively. Each group influences the system differently: $S_{t}^{(1)}$ affects only the reward, $S_{t}^{(2)}$ impacts both the action and the reward, and $S_{t}^{(3)}$ solely influences the action; see Figure E.1(a) for an illustration. As formally proven in Lemma E.1 (see Appendix E.1):
  • The type-I MSA selects the first two groups $S_{t}^{(1)}$ and $S_{t}^{(2)}$;
  • The proposed type-II MSA selects the last two groups $S_{t}^{(2)}$ and $S_{t}^{(3)}$;
Example II (An MDP example): We next consider an MDP with three groups of states, depicted in Figure E.1(b). Key observations from this example are as follows: (i) The reward depends on the state only through the first group of variables; (ii) The evolution of the state depends only on the second group. Specifically, the second group evolves first at each time step and subsequently influences the rest two groups; (iii) The behavior policy depends only on the last group; (iv) The second group is directly influenced by the previous action. Based on these observations, we show have that:
  • According to (ii) and (iv), selecting the second group achieves both transition-irrelevance (see the second equation in 5) and backward-transition-irrelevance (see (7)).
Consequently, the type-I and type-II MSAs select the first two and last two groups of variables, respectively. The iterative procedure again selects their intersection $S_{t}^{(2)}$ and converges in two iterations, leading to in a smaller state space. The reader is referred to Appendix E.1 for formal justifications.
In both examples, we have demonstrated the advantage of DSA in reducing state space cardinality over non-iterative procedures.
E.1 Lemma E.1 and its proof
We first state Lemma E.1.
Lemma E.1
@@ Appendix E Toy examples
  • The type-I MSA selects the first two groups $S_{t}^{(1)}$ and $S_{t}^{(2)}$;
  • The proposed type-II MSA selects the last two groups $S_{t}^{(2)}$ and $S_{t}^{(3)}$;
@@
We next prove this lemma. Notice that reward-irrelevance requires the reward function (i.e., the conditional mean of the immediate reward given the state-action pair) to depend on the state only through its abstraction. Under the deterministic reward assumption in Lemma E.1, such a conditional mean independence is equivalent to conditional independence. In other words, reward-irrelevance is achieved if the reward is conditionally independent of the state given the action and the abstract state.
E.1.1 Proof for the bandit example